# MokA: Multimodal Low-Rank Adaptation for MLLMs

### v1

@@ 4.2 Visual-text and speech-text scenarios
To further validate our method across a broader range of multimodal scenarios, we conducted experiments beyond the challenging audio-visual-text case. Specifically, our method is further verified on two representative multimodal scenarios: visual-text and speech-text. For these tasks, we adopted three different LLM backbones, LLaMA2, LLaMA3, and Qwen2. The corresponding results are presented inΒ TableΒ 2 andΒ TableΒ 3. The experimental results demonstrate that our method achieves stable and consistent performance gains across multiple benchmark datasets, further confirming its effectiveness. This indicates the versatility of MokA in handling different multimodal combinations and LLM architectures.
@@
| LoRAMoEΒ [7] | 27.63 | 33.97 | |||
| DoRAΒ [19] | 26.73 | 34.36 | 1 | 1 | |
| HydraLoRAΒ [27] | 29.13 | 31.66 | 1 | ||
| Uni-modal LoRAΒ [1] | 38.14 | 35.14 | |||
@@
| LoRAMoEΒ [7] | 50.75 | 42.99 | N | N | |
| DoRAΒ [19] | 52.55 | 44.11 | 1 | 1 | |
| HydraLoRAΒ [27] | 53.45 | 43.94 | 1 | ||
| Uni-modal LoRAΒ [1] | 54.05 | 47.01 | |||
@@ 4.3 Partial modality inference of MokA
To further examine how effectively MokA leverages tokens from different modalities, we also conduct partial modality inference experiments where only tokens from a selected modality are passed through the LoRA adaptation pathway at the first generation during inference. It should be noted that the evaluated model is MokA w/o cross-attention, as cross-attention computation requires the presence of both text and non-text tokens. The results, presented inΒ FigureΒ 5, show that MokA w/o cross-attention significantly enhances the utilization of individual modalities compared to LoRA (as shown inΒ FigureΒ 1d-1f). These findings highlight that the multimodal-aware design of MokA facilitates more effective use of all available modalities.
@@ 4.4 Cross-modal interaction variants
In the original MokA framework, cross-attention is employed to explicitly strengthen the interaction between text and non-text tokens, thereby facilitating improved cross-modal adaptation. As previously discussed, alternative modules that similarly enhance this interaction can also be considered. In this section, we explore several variants of the cross-modal interaction module, as summarized inΒ TableΒ 4. The cross-attention* variant also adopts a cross-attention mechanism; however, it uses text tokens as queries. Consequently, the updated text tokens integrate information from the relevant non-text tokensβreversing the direction of interaction compared to the original MokA. The naive interaction variant performs a simple, uniform mapping from text tokens to non-text tokens without employing any attention mechanism.
@@ 4.5 Ablation study
To thoroughly validate the efficacy of our method, we conduct ablation studies across all three multimodal scenarios. Results are shown inΒ TableΒ 5. Based on the results, even without the cross-attention module, MokA w/o CA outperforms the LoRA baseline, demonstrating the effectiveness of enhancing unimodal adaptation. Furthermore, the introduction of the cross-attention module leads to additional performance improvements, indicating the benefit of explicitly enhancing cross-modal adaptation. These results indicate the necessity of each part in MokA.
@@ 4.6 Efficiency evaluation
| LoRAΒ [11] | 1.27% | 1.00 | 70.28 | ||||||
| Multiple LoRA | 1.43% | 1.36 | 68.20 | ||||||
1.33% | 1.13 | 74.23 |
To enable a more comprehensive comparison, we further evaluate the proposed MokA and LoRA baselines on the proportion of trainable parameters in the full model, and inference latency. As reported inΒ TableΒ 6, although MokA introduces additional parameters due to the inclusion of more low-rank matrices, the increase is quite modest compared to the full LLM. Also, despite MokA incurs a slight increase in inference latency compared to standard LoRA, it achieves a notable performance gain of on the POPE benchmark. These results suggest that the additional computational cost introduced by MokA is acceptable, and the performance improvement is considerable.
@@ 5 Related works
MLLMs built upon powerful LLM backbones are increasingly demonstrating impressive capabilities across diverse downstream tasks. However, fine-tuning these models remains computationally expensive, prompting growing interest in parameter-efficient fine-tuning (PEFT) techniques that reduce memory and storage overhead during adaptation. Among them, LoRA has emerged as a widely adopted, and researchers have proposed several variants to further improve its efficiency and flexibilityΒ [7, 19, 27, 1]. For instance, LoRAMoEΒ [7] introduces multiple LoRA heads combined via a gating mechanism, while DoRAΒ [19] focuses solely on optimizing the gradient direction, enabling more efficient updates. Despite these advancements, most PEFT strategies for MLLMs are direct extensions of LLM techniques and fail to account for the inherent characteristics of multimodal learning. To address this gap, we propose MokA, a fine-tuning strategy specifically designed for MLLMs. It explicitly ensures both unimodal and cross-modal adaptation to better preserve unimodal representations and enhance cross-modal interaction, offering a targeted solution for efficient and effective multimodal adaptation.
@@ Appendix A Extension to Qwen3
To further validate our MokA method, we equip it with the latest Qwen3Β [32] model. Qwen3-8B is used as the LLM base model. Throughout the training process, weights of LLM are kept frozen.
Results are shown inΒ TableΒ 7. We conduct experiments in the audio-visual-text case. Our method can still deliver improvements compared to the LoRA baseline, with the latest Qwen3 model. These findings provide additional evidence of the effectiveness and scalability of MokA.
@@ Appendix B MokA with linear projection
The reason is that low-rank matrices of each modality actually can be considered as the linear projection in attention in this case. Therefore, we do not introduce other linear projections in the cross-attention module. In addition, projections of key and value are shared in this case (). In fact, this kind of projection sharing strategy has been widely used to increase the efficiency of attentionΒ [13, 30]. For example, in the LinformerΒ [30], it also utilizes the sharing key-value projection strategy to reduce computation cost. Whatβs more, here non-text tokens are used as queries to merge textual information into non-text modalities. In this way, cross-modal interaction is explicitly enhanced, while text tokens are kept unchanged to avoid potential disruption to the modelβs original strong text understanding capability.
@@
In Β TableΒ 8, we provide the results in the audio-visual-text and visual-text cases. Based on the results, MoKA with linear projection, yields improvements compared with LoRA baseline. However, it does not consistently outperform the original MoKA and introduces additional trainable parameters along with increased computational overhead.
@@
Appendix C Case study of cross-attention in MokA
In this section, we conduct a case study on the cross-attention module in MokA. This module explicitly integrates task description information from text tokens with non-text tokens, thereby facilitating cross-modal interaction. Here we provide two samples from an audio-visual-text scenario, as shown inΒ FigureΒ 6. The visualization shows the cross-attention weights of the at the th layer. The LLM model is LLaMA2. The results indicate that during the process of explicit cross-modal integration (e.g., cross-attention), text tokens that are more related to a given modality can receive higher attention weights. For example, the token βsoundβ receives greater attention in relation to the audio modality. This cross-modal integration can better facilitate the alignment and interaction between text tokens and non-text tokens.
Appendix D Broader impacts
@@ Appendix E Datasets
Video-LLaVAΒ [17] used a mixed dataset of images and videos for video captioning and image captioning tasks. The dataset includes a 665k image-text instruction and a 100k video-text instruction. This dataset is used for pre-training the visual branch.
AudioCapsΒ [12] dataset is used for the audio captioning task. It consists of 46K pairs of audio clips and text descriptions. This dataset is used for pre-training the audio branch.
GigaSpeech-MΒ [3] is a 1000h dataset for speech recognition task. This dataset is used for pre-training the speech branch.
MUSIC-AVQAΒ [15] is an audio-visual-text dataset, which is introduced to support spatio-temporal understanding of musical content. It offers 45K QA pairs across 33 question templates that span multiple modalities and question types.
AVEΒ [28] is an audio-visual-text dataset. It focuses on the audio-visual event localization task. This dataset covers 28 event classes and consists of 4,143 samples.
LLaVA-Instruct-150KΒ [18] is a set of GPT-generated multimodal instruction-following data. It is used for instruction fine-tuning for the visual-text case.
LibriSpeechΒ [21] is a 960-hour dataset. We use the instruction fromΒ [26] for instruction fine-tuning of the speech-text case.
MMEpercepΒ [8] is the perception subset of the MME benchmark, covering a total of 10 subtasks for the evaluation of the visual-text perception ability.
MMBenchΒ [20]is a collection of benchmarks to evaluate the visual-text understanding capability. It has 3,000 multiple-choice questions covering object detection, text recognition, action recognition, image captioning, relation reasoning, and so on.
POPEΒ [16] is a benchmark that is used for evaluating the visual-text understanding ability of MLLM. The used image is the test set of MSCOCO dataset.
SEED-BenchΒ [14] consists of 19K multiple-choice questions with accurate human annotations for evaluating the visual-text understanding ability of MLLM.
MMAUmin-speechΒ [25] is the speech subset of MMAU-mini benchmark. This benchmark is used for evaluating the speech-text understanding ability of MLLM.
AIR-Benchspeech-enΒ [34] is the English speech subset of the foundation part of AIR-Bench. This benchmark is used for evaluating the speech-text understanding ability of MLLM.

### v2

@@ 4.2 Visual-text and speech-text scenarios
To further validate our method across a broader range of multimodal scenarios, we conducted experiments beyond the challenging audio-visual-text case. Specifically, our method is further verified on two representative multimodal scenarios: visual-text and speech-text. For these tasks, we adopted three different LLM backbones, LLaMA2, LLaMA3, and Qwen2. The corresponding results are presented in Table 2 and Table 3. The experimental results demonstrate that our method achieves stable and consistent performance gains across multiple benchmark datasets, further confirming its effectiveness. This indicates the versatility of MokA in handling different multimodal combinations and LLM architectures.
@@
| LoRAMoE [dou2023loramoe] | 27.63 | 33.97 | |||
| DoRA [liu2024dora] | 26.73 | 34.36 | 1 | 1 | |
| HydraLoRA [tian2024hydralora] | 29.13 | 31.66 | 1 | ||
| Uni-modal LoRA [abouelenin2025phi] | 38.14 | 35.14 | |||
@@
| LoRAMoE [dou2023loramoe] | 50.75 | 42.99 | N | N | |
| DoRA [liu2024dora] | 52.55 | 44.11 | 1 | 1 | |
| HydraLoRA [tian2024hydralora] | 53.45 | 43.94 | 1 | ||
| Uni-modal LoRA [abouelenin2025phi] | 54.05 | 47.01 | |||
@@ 4.3 Partial modality inference of MokA
To further examine how effectively MokA leverages tokens from different modalities, we also conduct partial modality inference experiments where only tokens from a selected modality are passed through the LoRA adaptation pathway at the first generation during inference. It should be noted that the evaluated model is MokA w/o cross-attention, as cross-attention computation requires the presence of both text and non-text tokens. The results, presented in Figure 5, show that MokA w/o cross-attention significantly enhances the utilization of individual modalities compared to LoRA (as shown in Figure 1d-1f). These findings highlight that the multimodal-aware design of MokA facilitates more effective use of all available modalities.
@@ 4.4 Cross-modal interaction variants
In the original MokA framework, cross-attention is employed to explicitly strengthen the interaction between text and non-text tokens, thereby facilitating improved cross-modal adaptation. As previously discussed, alternative modules that similarly enhance this interaction can also be considered. In this section, we explore several variants of the cross-modal interaction module, as summarized in Table 4. The cross-attention* variant also adopts a cross-attention mechanism; however, it uses text tokens as queries. Consequently, the updated text tokens integrate information from the relevant non-text tokens—reversing the direction of interaction compared to the original MokA. The naive interaction variant performs a simple, uniform mapping from text tokens to non-text tokens without employing any attention mechanism.
@@ 4.5 Ablation study
To thoroughly validate the efficacy of our method, we conduct ablation studies across all three multimodal scenarios. Results are shown in Table 5. Based on the results, even without the cross-attention module, MokA w/o CA outperforms the LoRA baseline, demonstrating the effectiveness of enhancing unimodal adaptation. Furthermore, the introduction of the cross-attention module leads to additional performance improvements, indicating the benefit of explicitly enhancing cross-modal adaptation. These results indicate the necessity of each part in MokA.
@@ 4.6 Efficiency evaluation
| LoRA [hu2022lora] | 1.27% | 1.000 | 70.28 | ||||||
| Multiple LoRA | 1.43% | 1.006 | 68.20 | ||||||
1.33% | 1.069 | 74.23 |
To enable a more comprehensive comparison, we further evaluate the proposed MokA and LoRA baselines on the proportion of trainable parameters in the full model, and inference latency. As reported in Table 6, although MokA introduces additional parameters due to the inclusion of more low-rank matrices, the increase is quite modest compared to the full LLM. Also, despite MokA incurs a slight increase in inference latency compared to standard LoRA, it achieves a notable performance gain of on the POPE benchmark. These results suggest that the additional computational cost introduced by MokA is acceptable, and the performance improvement is considerable. More detailed efficiency evaluations are provided in Appendix D.
@@ 5 Related works
MLLMs built upon powerful LLM backbones are increasingly demonstrating impressive capabilities across diverse downstream tasks [wang2023large, jiang2025cmsl]. However, fine-tuning these models remains computationally expensive, prompting growing interest in parameter-efficient fine-tuning (PEFT) techniques that reduce memory and storage overhead during adaptation. Among them, LoRA has emerged as a widely adopted, and researchers have proposed several variants to further improve its efficiency and flexibility [dou2023loramoe, liu2024dora, tian2024hydralora, abouelenin2025phi]. For instance, LoRAMoE [dou2023loramoe] introduces multiple LoRA heads combined via a gating mechanism, while DoRA [liu2024dora] focuses solely on optimizing the gradient direction, enabling more efficient updates. Despite these advancements, most PEFT strategies for MLLMs are direct extensions of LLM techniques and fail to account for the inherent characteristics of multimodal learning. To address this gap, we propose MokA, a fine-tuning strategy specifically designed for MLLMs. It explicitly ensures both unimodal and cross-modal adaptation to better preserve unimodal representations and enhance cross-modal interaction, offering a targeted solution for efficient and effective multimodal adaptation.
@@ 6 Discussion
Acknowledgement
The project was supported by the fund for building world-class universities (disciplines) of Renmin University of China, sponsored by CCF-Zhipu.AI Large Model Innovation Fund, and National Natural Science Foundation of China (NO.62106272). The project was also supported by the Shanghai Municipal Science and Technology Major Project, and Shanghai Artificial Intelligence Laboratory.
@@ Appendix A Extension to Qwen3
To further validate our MokA method, we equip it with the latest Qwen3 [yang2025qwen3technicalreport] model. Qwen3-8B is used as the LLM base model. Throughout the training process, weights of LLM are kept frozen.
Results are shown in Table 7. We conduct experiments in the audio-visual-text case. Our method can still deliver improvements compared to the LoRA baseline, with the latest Qwen3 model. These findings provide additional evidence of the effectiveness and scalability of MokA.
@@ Appendix B MokA with linear projection
The reason is that low-rank matrices of each modality actually can be considered as the linear projection in attention in this case. Therefore, we do not introduce other linear projections in the cross-attention module. In addition, projections of key and value are shared in this case (). In fact, this kind of projection sharing strategy has been widely used to increase the efficiency of attention [kitaev2020reformer, wang2020linformer]. For example, in the Linformer [wang2020linformer], it also utilizes the sharing key-value projection strategy to reduce computation cost. What’s more, here non-text tokens are used as queries to merge textual information into non-text modalities. In this way, cross-modal interaction is explicitly enhanced, while text tokens are kept unchanged to avoid potential disruption to the model’s original strong text understanding capability.
@@
In Table 11, we provide the results in the audio-visual-text and visual-text cases. Based on the results, MoKA with linear projection, yields improvements compared with LoRA baseline. However, it does not consistently outperform the original MoKA and introduces additional trainable parameters along with increased computational overhead.
Appendix C More than three modality case
In this section, we extend our experiments to scenarios with more than three modalities. Specifically, we consider a four-modality setting involving audio, visual, point cloud, and language data, and evaluated it on the MCUB-3 benchmark [chen2024model]. In this 3+ modality case, LoRA has an accuracy of 37.41%. Our MokA has an accuracy of 45.58%, indicating its scalability and effectiveness under 3+ modality cases.
Appendix D Efficiency evaluation
Compared to LoRA, MokA introduces additional A matrices and a cross-attention module. However, it should be noted that the additional computational cost of MokA only comes from the cross-attention module, which is flexible and can be replaced by a more efficient strategy if needed. In this section, we provide a more detailed efficiency analysis comparing FLOPs, GPU memory usage, and average forward time per sample (proportional to training time). For clarity, we report these metrics for the two-modality (VL), three-modality (AVL), and four-modality (AVPL) settings. As shown in the tables, MokA’s extra computational and memory cost remains limited and acceptable for typical multimodal scenarios.
| VL (MME_) | FLOPs | Memory Usage | Avg. forward time/sample |
| LoRA | 1.000x | 1.000x | 1.000x |
1.009x | 1.001x | 1.069x |
| AVL (MUSIC-AVQA) | FLOPs | Memory Usage | Avg. forward time/sample |
| LoRA | 1.000x | 1.000x | 1.000x |
1.021x | 1.001x | 1.134x |
| AVPL (MCUB-3) | FLOPs | Memory Usage | Avg. forward time/sample |
| LoRA | 1.000x | 1.000x | 1.000x |
1.013x | 1.002x | 1.213x |
Appendix E Audio-visual interaction
In the original MokA, only the attention between text and non-text tokens is considered. It is motivated by the fact that text modality typically conveys the question or task description, while the audio and visual modalities provide environmental information, in the instruction. Therefore, cross-attention is applied to explicitly enhance the interaction between the task (text token) and environment (non-text token). It is also worth exploring whether further interactions between the scene modalities themselves—i.e., audio and visual—can be beneficial. To this end, we conduct additional ablation studies. Experiments are conducted based on LLaMA2. The table reports the results for audio-visual attention with both audio as query and video as query. The results show that introducing additional audio-visual attention can bring gains, but it is not very noticeable. The potential benefit from further enhancing explicit audio-visual interactions is relatively limited.
| Method | MUSIC-AVQA | AVE |
| LoRA | 73.41 | 69.84 |
75.71 | 74.68 |
| MokA w/ audio query att. | 75.78 | 74.53 |
| MokA w/video query att. | 75.76 | 74.81 |
Appendix F Ablation study of rank
In this section, we conduct experiments of MokA with different ranks. It consistently outperforms the LoRA baseline across different rank settings.
| Method | MUSIC-AVQA | AVE | Rank |
| LoRA | 73.41 | 69.84 | r=4 |
| LoRA | 73.56 | 70.01 | r=8 |
| LoRA | 73.73 | 70.07 | r=12 |
75.71 | 74.68 | r=4 |
74.68 | 74.71 | r=8 |
74.89 | 74.36 | r=12 |
Appendix G Case study of cross-attention in MokA
In this section, we conduct a case study on the cross-attention module in MokA. This module explicitly integrates task description information from text tokens with non-text tokens, thereby facilitating cross-modal interaction. Here we provide two samples from an audio-visual-text scenario, as shown in Figure 6. The visualization shows the cross-attention weights of the at the th layer. The LLM model is LLaMA2. The results indicate that during the process of explicit cross-modal integration (e.g., cross-attention), text tokens that are more related to a given modality can receive higher attention weights. For example, the token “sound” receives greater attention in relation to the audio modality. This cross-modal integration can better facilitate the alignment and interaction between text tokens and non-text tokens.
@@ Appendix G Case study of cross-attention in MokA
Appendix H Broader impacts
@@ Appendix I Datasets
Video-LLaVA [lin2023video] used a mixed dataset of images and videos for video captioning and image captioning tasks. The dataset includes a 665k image-text instruction and a 100k video-text instruction. This dataset is used for pre-training the visual branch.
AudioCaps [kim2019audiocaps] dataset is used for the audio captioning task. It consists of 46K pairs of audio clips and text descriptions. This dataset is used for pre-training the audio branch.
GigaSpeech-M [chen2021gigaspeech] is a 1000h dataset for speech recognition task. This dataset is used for pre-training the speech branch.
MUSIC-AVQA [li2022learning] is an audio-visual-text dataset, which is introduced to support spatio-temporal understanding of musical content. It offers 45K QA pairs across 33 question templates that span multiple modalities and question types.
AVE [tian2018audio] is an audio-visual-text dataset. It focuses on the audio-visual event localization task. This dataset covers 28 event classes and consists of 4,143 samples.
LLaVA-Instruct-150K [liu2023visual] is a set of GPT-generated multimodal instruction-following data. It is used for instruction fine-tuning for the visual-text case.
LibriSpeech [panayotov2015librispeech] is a 960-hour dataset. We use the instruction from [tang2023salmonn] for instruction fine-tuning of the speech-text case.
MMEpercep [fu2024mmecomprehensiveevaluationbenchmark] is the perception subset of the MME benchmark, covering a total of 10 subtasks for the evaluation of the visual-text perception ability.
MMBench [liu2024mmbench]is a collection of benchmarks to evaluate the visual-text understanding capability. It has 3,000 multiple-choice questions covering object detection, text recognition, action recognition, image captioning, relation reasoning, and so on.
POPE [li2023evaluating] is a benchmark that is used for evaluating the visual-text understanding ability of MLLM. The used image is the test set of MSCOCO dataset.
SEED-Bench [li2023seedbenchbenchmarkingmultimodalllms] consists of 19K multiple-choice questions with accurate human annotations for evaluating the visual-text understanding ability of MLLM.
MMAUmini-speech [sakshi2024mmau] is the speech subset of MMAU-mini benchmark. This benchmark is used for evaluating the speech-text understanding ability of MLLM.
AIR-Benchspeech-en [yang2024air] is the English speech subset of the foundation part of AIR-Bench. This benchmark is used for evaluating the speech-text understanding ability of MLLM.